### v1

# TimeBlind: A Spatio-Temporal Compositionality Benchmark for Video LLMs

Abstract
Fine-grained spatio-temporal understanding is essential for video reasoning and embodied AI. Yet, while Multimodal Large Language Models (MLLMs) master static semantics, their grasp of temporal dynamics remains brittle. We present TimeBlind, a diagnostic benchmark for compositional spatio-temporal understanding. Inspired by cognitive science, TimeBlind categorizes fine-grained temporal understanding into three levels: recognizing atomic events, characterizing event properties, and reasoning about event interdependencies. Unlike benchmarks that conflate recognition with temporal reasoning, TimeBlind leverages a minimal-pairs paradigm: video pairs share identical static visual content but differ solely in temporal structure, utilizing complementary questions to neutralize language priors. Evaluating over 20 state-of-the-art MLLMs (e.g., GPT-5, Gemini 3 Pro) on 600 curated instances (2400 video-question pairs), reveals that the Instance Accuracy (correctly distinguishing both videos in a pair) of the best performing MLLM is only 48.2%, far below the human performance (98.2%). These results demonstrate that even frontier models rely heavily on static visual shortcuts rather than genuine temporal logic, positioning TimeBlind as a vital diagnostic tool for next-generation video understanding. Dataset and code are available at https://baiqi-li.github.io/timeblind_project/.
1 Introduction
Figure 1: An example video pair that shares identical static visual content but differs solely in motion dynamics. The top video shows a person shaking a cup while making coffee, while the bottom video shows them holding it still. Even the most advanced models like GPT-5 and Gemini 3 Pro fail to distinguish the actions in the video pair.
Table 1: Comparison with prior temporal video benchmarks. We compare coverage of temporal reasoning categories and anti-shortcut design features. TimeBlind uniquely covers all 13 Allen temporal relations, includes causal and comparative reasoning, and employs both paired videos and complementary questions to eliminate static and linguistic shortcuts. The rightmost column shows the gap between human and best model performance, indicating benchmark difficulty.
Temporal Reasoning Coverage Anti-Shortcut Design
Benchmark Size Event Types Event Attributes Temporal Topologies Causal Cross-Event Comparison Video Pairs Compl. Questions Human–Model Gap
MVBench (li2024mvbench) 4K 2 2 2
TOMATO (shangguan2024tomato) 1.4K 2 4 2 57.3%
Vinoground (zhang2024vinoground) 2K 2 1 2 40.0%
TempCompass (liu2024tempcompass) 4K 2 2 2
TimeBlind (Ours) 2.4K 2 6 13 50.0%
Covers all 13 event relations in Allen’s Interval Algebra (allen1983maintaining): before, after, meets, met-by, overlaps, overlapped-by, starts, started-by, finishes, finished-by, during, contains, equals.
Fine-grained spatio-temporal understanding is fundamental for long-horizon video reasoning (wang2025lvbench; pan2025basket) and embodied AI (tang2025video; yang2025cambrian). Beyond simply recognizing what is present, an intelligent system must infer what changes, how it changes, and how multiple changes compose into causal structures. While recent Multimodal Large Language Models (MLLMs) (gemini; openai2025gpt5systemcard; bai2025qwen3vltechnicalreport; clark2025molmo2; cho2025perceptionlm) demonstrate impressive performance on general benchmarks, their “sense of time” remains surprisingly brittle. As shown in Figure 1, even frontier models (e.g., GPT-5 and Gemini 3 Pro) struggle to distinguish atomic actions (e.g., shaking vs. holding) in videos as short as 10 seconds. Moreover, models often misjudge relative dynamics (e.g., accelerating vs. decelerating), and frequently fail to resolve linguistic temporal connectives (e.g., “as soon as”). This discrepancy suggests that current benchmark scores (jang2017tgif; activitynet; li2024mvbench; zhou2025mlvu; fu2025video; hu2025video) severely overestimate genuine temporal understanding capabilities.
The root issue lies in evaluation design. Existing benchmarks (activitynet; li2024mvbench; liu2024tempcompass; fu2025video) rarely isolate temporal structure as the sole discriminative factor. As a result, many models exploit “static shortcuts”—correlating visual entities with answers without modeling time (lei2023revealing; krojer2025shortcut). Furthermore, language priors (li2024naturalbench) allow models to guess answers based on textual plausibility. To truly diagnose understanding, an evaluation must hold the visual content constant and vary only the temporal dynamics.
To address this, we introduce TimeBlind, a diagnostic benchmark for compositional spatio-temporal understanding (Table 1). TimeBlind adopts a minimal-pairs design: each instance contains two videos with near-identical static visual content that differ solely in their temporal structure. To eliminate language priors, we utilize complementary questions where the correct answer flips between the paired videos. By removing static and linguistic shortcuts, TimeBlind forces models to rely exclusively on temporal evidence. Unlike large-scale noisy benchmarks, TimeBlind prioritizes high-fidelity diagnostic precision. Similar to Winoground (thrush2022winoground), each instance serves as a rigorous test for a specific cognitive primitive, where high-quality annotations are prioritized over scale.
Crucially, unlike prior work (zhang2024vinoground; krojer2025shortcut), TimeBlind evaluates logical composition, not just perception. Inspired by cognitive science, we extend the theory of Image Compositionality (krishna2017visual; li2024genai) to the temporal domain, organizing the benchmark around a hierarchical taxonomy: (i) Events, addressing what happened (e.g., changes in object attributes or action understanding); (ii) Event Attributes, describing how events unfold (e.g., speed, magnitude of change); and (iii) Structural Event Logic, examining how multiple events are composed (e.g., temporal topology, causality, and cross-event comparison). Within this framework, we encompass a diverse set of 11 fine-grained categories in real-world scenarios. For example, regarding Temporal Topology in Structural Event Logic, we incorporate Allen’s Interval Algebra (allen1983maintaining) by constructing videos that cover all 13 temporal event relations (e.g., overlaps, meets, equals) and designing questions that distinguish specific topology differences within each pair, going beyond simple sequencing relations (e.g., before, after) used in prior work.
We evaluate over 20 state-of-the-art MLLMs, including leading proprietary models (e.g., Gemini 3 Pro (gemini), GPT-5 (openai2025gpt5systemcard)) as well as strong open-source models (e.g., Qwen3-VL (bai2025qwen3vltechnicalreport), Molmo2 (clark2025molmo2), PLM (cho2025perceptionlm)), on 600 curated instances (totaling 2,400 video-question pairs). We find that even the top-performing model, Gemini-3 Pro, achieves only 48.2% Instance Accuracy (I-Acc), which requires correctly distinguishing both videos in a pair, falling far below human performance (98.2%). Furthermore, while advanced models like Gemini-3 Pro and GPT-5 demonstrate reasonable proficiency in recognizing isolated events (achieving 49.2% and 58.3% I-Acc, respectively), their performance degrades sharply on event attributes—such as Speed (slowly vs. rapidly), and Force (forcefully vs. gently)—dropping to 36.7% and 32.3% I-Acc. Moreover, extensive ablation studies demonstrate that model performance remains poor even with increased input frames or test-time reasoning, as GPT-5 gains only 3.3% I-Acc, indicating models still depend mostly on visual and linguistic shortcuts rather than temporal reasoning. We believe TimeBlind will serve as a vital diagnostic tool for evaluating and developing MLLMs capable of genuine temporal logic.
In summary, our main contributions are:
  • TimeBlind Benchmark. A diagnostic, minimal-pairs benchmark that isolates temporal structure while minimizing static shortcuts and language priors.
  • Taxonomy of Temporal Compositionality. A cognitive hierarchy—Events, Attributes, and Structural Logic—that guides systematic benchmark construction and evaluation.
  • Diagnostic Findings. An evaluation of over 20 SOTA MLLMs revealing a significant gap between perceived and actual temporal reasoning capabilities.
2 Related Work
Figure 2: TimeBlind Taxonomy and Statistics. Left: We structure the evaluation into 11 fine-grained spatio-temporal compositional categories spanning three high-level aspects: Atomic Events (what changes), Parametric Event Attributes (how it changes), and Structural Event Logic (how events compose). Top Right: Distribution of video lengths across the benchmark, showing that most videos fall within the 0–15 seconds. Bottom Right: Distribution of question word counts, indicating that most questions are under 30 words. Overall, our benchmark features a structured taxonomy with diverse categories while maintaining short videos and concise questions.
VideoQA Benchmarks. Early VideoQA datasets (activitynet; tgif; perception; xiao2021next; lei2018tvqa; xu2016msr; wu2024star) focus on simple scenarios with short clips and limited question types. Recent benchmarks address comprehensive evaluation (fu2025video; li2024mvbench; ma2025videoeval), complex reasoning (nagrani2025minerva; cheng2025video; hu2025video; song2025video), long-form understanding (mangalam2023egoschema; wang2025lvbench; wu2024longvideobench; song2024moviechat; zhou2025mlvu; yang2025egolife), and domain-specific settings (pan2025basket; long2025adsqa; yi2025exact). However, most do not isolate temporal structure as the sole discriminative factor, allowing models to exploit static shortcuts—relying on object co-occurrence or language priors without genuinely modeling temporal dynamics (lei2023revealing; krojer2025shortcut; li2024naturalbench).
Evaluating Spatio-Temporal Compositionality. To assess the spatio-temporal understanding capabilities of MLLMs, several recent benchmarks have introduced diverse aspects of visual reasoning. In the image understanding domain, BLINK (fu2024blink) reformats and groups classic vision problems into perception-centric multiple-choice questions and carefully removes language priors. In the video understanding domain, several works explicitly target temporal understanding by designing temporally challenging questions (johnson2017clevr; shangguan2024tomato; cai2024temporalbench; xue2025seeing). Beyond single-video evaluation, a growing line of work adopts paired video–question protocols to more precisely diagnose temporal reasoning. TempCompass (liu2024tempcompass) constructs paired videos by systematically manipulating an original video, such as reversing playback or altering temporal speed, thereby isolating temporal understanding from “static shortcuts.” VinoGround (zhang2024vinoground) further strengthens this paradigm by requiring models to answer identical questions over paired videos, where the correct answer is determined solely by temporal differences, effectively removing language priors. Follow-up works like GLIMPSE (zhou2025glimpse) and MVP (krojer2025shortcut) scale this approach to test physical and visual-centric reasoning. Different from prior works, TimeBlind achieves spatio-temporal compositionality through a carefully curated formal taxonomy. Drawing inspiration from cognitive event perception (bach1986algebra) and extending the theory of image compositionality (krishna2017visual; thrush2022winoground; li2024naturalbench), we decompose temporal reasoning into atomic Events, parametric Event Attributes, and Structural Logic.
3 The TimeBlind Benchmark
Figure 3: Overview of the TimeBlind data construction pipeline. Stage 1 (Schema Generation): We prompt GPT-5 to generate paired complementary questions targeting temporal differences. Stage 2 (Video Acquisition): We collect one video pair that matches the generated schema from one of the following sources: (i) Retrieving videos from the internet, (ii) Recording videos with humans, or (iii) Generating videos via simulation (e.g., Unity). We then pair these videos with the questions to form a candidate TimeBlind instance. Stage 3 (Manual Review): Human annotators manually review each instance to ensure: (i) Static Consistency (videos share identical static content), (ii) Temporal Minimality (the pair differs only in the targeted temporal factor), and (iii) Question Validity (QA pairs are clear and correct).
In this section, we describe the design of TimeBlind, a diagnostic benchmark for rigorously evaluating compositional spatio-temporal understanding. Our design is grounded in two principles: (i) a temporal minimal-pair protocol that minimizes visual shortcuts and language priors, and (ii) a cognition-inspired compositionality taxonomy that decomposes temporal reasoning into three primitives: Events, Event Attributes, and Structural Event Logic.
3.1 Task Formulation
We structure each instance as $\mathcal{T}=(v_{1},v_{2},q_{1},q_{2})$, following the design below. This design creates a discriminative challenge where shortcut solutions are effectively “canceled out,” forcing the model to rely on temporal evidence.
Temporal Minimal Pairs. The video pair $(v_{1},v_{2})$ constitutes a minimal pair: the videos share identical static content (e.g., objects, background) but differ along an isolated temporal axis defined in our taxonomy (Section 3.2). Consequently, a model cannot distinguish $v_{1}$ from $v_{2}$ without explicitly modeling the temporal features (e.g., opening vs. closing).
Logical Complementarity. The questions $(q_{1},q_{2})$ are complementary: for any question, the ground-truth answer flips between the two videos (e.g., $Ans(v_{1},q_{j})\neq Ans(v_{2},q_{j})$ for $j\in\{1,2\}$). This design neutralizes language priors, preventing models from exploiting textual plausibility.
Diagnostic Metrics. Following rigorous diagnostic protocols (thrush2022winoground; li2024naturalbench), we report a hierarchy of metrics to mitigate shortcuts and assess the models’ genuine temporal understanding:
  • Accuracy (Acc): Standard accuracy computed over all individual video–question trials.
  • Video Accuracy (V-Acc): Measures visual consistency; correct only if the model answers both questions correctly for a single video $v_{j}$.
  • Question Accuracy (Q-Acc): Measures textual consistency; correct only if the model answers question $q_{i}$ correctly for both videos.
  • Instance Accuracy (I-Acc): An instance is correct if and only if the model solves all four trials. I-Acc is our primary proxy for spatio-temporal understanding, as it necessitates reliably distinguishing the temporal difference between the paired videos.
3.2 A Hierarchy of Temporal Composition
To systematically evaluate temporal understanding, we organize TimeBlind around a compositional taxonomy mirroring a bottom-up cognitive process: from recognizing atomic changes (Events), to measuring their intrinsic properties (Event Attributes), and finally reasoning about their inter-dependencies (Structural Event Logic), shown in Figure 2.
I. Events. This foundational primitive tests the detection of atomic visual changes, demanding that models distinguish actual temporal evolution from static object presence.
  • Fine-Grained Actions: Discriminating actions sharing visual context but differing in temporal dynamics (e.g., opening vs. closing) or semantic nuance (e.g., placing vs. pushing).
  • State Transitions: Tracking object property changes over time (shape, size, color, emotion)—e.g., liquid changing from clear to opaque.
II. Event Attributes. This primitive tests the perception of how an event unfolds, requiring sensitivity to continuous or qualitative parameters rather than categorical recognition.
  • Kinematics: Properties governing motion in space-time, including Speed (fast/slow), Direction (forward/backward), Duration (brief/long), and Repetition (once/twice).
  • Dynamics & Manner: Properties reflecting physical forces or execution style, including Force (gentle/forceful) and Magnitude (slight/broad movement).
III. Structural Event Logic. This level tests how multiple events relate and compose into higher-order structures.
  • Temporal Topology: We adopt Allen’s Interval Algebra (allen1983maintaining) to comprehensively evaluate interval relations. Unlike prior work limited to simple sequencing (before/after), we cover all 13 relations in our videos, such as overlaps, meets, starts, and during.
  • Causal Contingency: Identifying causal links between events (e.g., distinguishing causation from correlation).
  • Cross-Event Comparison: Comparative reasoning across distinct events (e.g., “Does the person hold the cup longer than the book?”).
3.3 Data Construction Pipeline
We use a three-stage pipeline to collect high-quality TimeBlind data, as shown in Figure 3.
Step 1: Schema Generation. We prompt frontier LLMs (e.g., GPT-5) with the taxonomy definitions in Section 3.2 to generate structured specifications. Each output includes a pair of video descriptions (e.g., “a girl walks in a park and gradually speeds up” and “a girl walks in a park and gradually slows down”) that differ only in the target temporal factor (e.g., “Speed”), along with complementary questions (e.g., “Is the girl walking faster and faster?” and “Is the girl walking slower and slower?”).
Step 2: Video Acquisition. We collect video pairs from three sources: (i) internet retrieval (extracting temporally distinct segments from the same source video), (ii) human recording, or (iii) simulation (e.g., Unity) for precise temporal control. We then pair the collected videos with the questions to form a candidate TimeBlind instance.
Step 3: Rigorous Human Verification. Every instance undergoes strict manual review. Annotators verify (i) Static Consistency, ensuring both videos share identical static content; (ii) Temporal Minimality, confirming the pair differs only in the targeted temporal factor; and (iii) Question Validity, ensuring the QA pairs are clear and correct.
Statistics. TimeBlind comprises 2,400 curated video-question pairs, consisting of binary and two-choice multiple-choice questions. The video sources are distributed as follows: 24.0% from Internet Retrieval, 57.7% from Human Recording, and 18.3% from Simulation. Regarding category distribution, we maintain an approximate 1:1:1 ratio across Event, Event Attribute, and Structural Event Logic. Furthermore, within the Temporal Topology sub-category, we ensure a balanced distribution across the 13 distinct Allen interval types.
4 Experimental Setup
Baseline Models. To thoroughly assess the challenges posed by TimeBlind, we evaluate over 20 MLLMs, including both frontier proprietary models (e.g., GPT-5 (openai2025gpt5systemcard), Gemini 3 Pro (gemini), Claude (anthropic_claude_sonnet_4_5_2025), and Qwen3-VL Plus (bai2025qwen3vltechnicalreport)) and state-of-the-art open-source models (e.g., Qwen3-VL (bai2025qwen3vltechnicalreport), InternVL 3.5 (internvl3), Molmo2 (clark2025molmo2), PLM (cho2025perceptionlm), Keye-VL (team2025kwai), MiniCPM-V-4.5 (yu2025minicpm), GLM-4.1V (hong2025glm), and Eagle 2.5 (chen2025eagle2.5)).
Evaluation Metrics. We adopt the hierarchy of metrics defined in Section 3.1, including Standard Accuracy (Acc), Video Accuracy (V-Acc), Question Accuracy (Q-Acc), and our primary metric, Instance Accuracy (I-Acc).
Implementation Details. We conduct most experiments on 4 NVIDIA H100 GPUs, each with 96GB of memory. For proprietary models (e.g., GPT-5, Gemini 3 Pro) and large-scale open-source models (e.g., Qwen3-VL-235B), we utilize the official APIs. Unless otherwise specified, we sample videos uniformly at 1 FPS (the default frame rate for most models) and evaluate models in a zero-shot setting.
5 Experimental Results
5.1 Main Results
Table 2: Main Results on TimeBlind. We use uniform sampling at 1 FPS and evaluate all models using default configurations. Metrics are reported following Section 3, with I-Acc as the primary metric. The table is divided into Open-Source (grouped by size: <10<10B and >10>10B), and Closed-Source models. Our results suggest that all models perform poorly on fine-grained temporal video understanding, with none of the methods achieving over 50% I-Acc. Best results are shown in bold.
Model Size Q-Acc (%) V-Acc (%) Acc (%) I-Acc (%) $\uparrow$
Random Chance - 25.0 25.0 50.0 6.3
Human Evaluation - 98.8 98.9 99.3 98.2
[…]
Model Size $\leq 10$B
Keye-VL-1.5 (team2025kwai) 8B 17.5 26.5 55.6 6.6
LLaVA-Video (zhang2024video) 7B 18.1 34.8 57.8 9.6
InternVL3.5 (internvl3) 8B 30.9 39.0 59.3 13.3
PLM-8B (cho2025perceptionlm) 8B 24.4 32.3 60.6 13.9
Qwen3-VL (bai2025qwen3vltechnicalreport) 8B 34.5 43.6 64.8 19.7
MiniCPM-V-4.5 (yu2025minicpm) 8B 31.7 43.6 62.7 18.3
GLM-4.1V (hong2025glm) 9B 32.3 45.2 62.0 19.7
Qwen3-VL (bai2025qwen3vltechnicalreport) 4B 29.9 43.7 62.6 17.7
VideoChat-Flash (li2024videochat) 7B 31.7 45.4 64.9 19.7
Eagle2.5 (chen2025eagle2.5) 8B 32.2 45.7 64.7 20.2
Molmo2 (clark2025molmo2) 8B 41.0 52.4 68.7 31.2
Model Size $\geq 30$B
InternVL 3.5 (internvl3) 30B 31.4 41.3 61.1 17.0
LLaVA-Video (zhang2024video) 72B 28.8 45.0 63.2 18.5
Qwen3-VL (bai2025qwen3vltechnicalreport) 235B 38.1 51.7 66.9 25.8
[…]
Claude Sonnet 4.5 (anthropic_claude_sonnet_4_5_2025) - 30.6 38.5 59.2 13.6
Qwen3-VL Plus (bai2025qwen3vltechnicalreport) - 38.7 52.6 67.3 26.0
GPT-5 mini (openai2025gpt5systemcard) - 42.3 53.7 68.3 30.0
Gemini 2.5 Flash (google_gemini_2_5_thinking_2025) - 47.6 53.7 68.9 33.2
Gemini 2.5 Pro (google_gemini_2_5_thinking_2025) - 56.0 62.6 74.2 43.5
GPT-5 (openai2025gpt5systemcard) - 58.5 67.2 77.3 46.3
Gemini 3 Pro (gemini) - 60.2 66.0 76.2 48.2
In Table 5.1, we present the overall model performances on our benchmark, revealing that all models perform poorly in fine-grained temporal video understanding. While many models achieve relatively high standard accuracy (Acc), they struggle to achieve good Instance Accuracy (I-Acc), which requires correctly distinguishing both videos in a pair. For instance, despite leading models like GPT-5 and Gemini 3 Pro achieving high Acc scores of 77.3% and 76.2% respectively, they only reach 46.3% and 48.2% on I-Acc. This discrepancy indicates that models remain weak in temporal understanding and suggests that high performance on Acc is often driven by shortcuts, rather than a true understanding of temporal dynamics. Another interesting finding is that Question Accuracy (Q-Acc) is consistently lower than Video Accuracy (V-Acc) across all evaluated models. This indicates that models are more prone to hallucinating answers based on textual patterns than misinterpreting visual cues. From the results, we also observe that among open-source models, Molmo2-8B—despite its smaller size—outperforms a range of sub-10B models and even surpasses the much larger Qwen3-VL-235B by 5.4% on I-Acc, becoming the leading open-source model. This suggests that with effective design choices in model architecture, data curation, and training process, even smaller models can capture fine-grained temporal dynamics. Lastly, our results show a significant gap between open-source and proprietary models in understanding fine-grained temporal dynamics. Specifically, Molmo2-8B still lags substantially behind GPT-5 and Gemini 3 Pro, with gaps of 15.1% and 17.0% I-Acc, respectively.
Human Evaluation. We validate TimeBlind with four independent annotators (Table 5.1). Each annotator saw only one question and one video at a time, with the four $(v,q)$ pairs from each instance distributed across different annotators. Humans achieved 98.2% I-Acc—exceeding Gemini 3 Pro by 50%—showing that temporal dynamics in TimeBlind are clear to humans while remaining challenging for MLLMs.
5.2 Category-Wise Diagnosis
To pinpoint specific cognitive deficits, we analyze performance across the TimeBlind taxonomy in Table 5.2.
  • Performance Across Hierarchical Levels. We observe notable performance gaps across different temporal understanding tasks, with models generally performing better on discrete Events (e.g., distinguishing atomic actions) compared to continuous Event Attributes or Structural Event Logic. For example, GPT-5 achieves 58.3% accuracy in the Event category (peaking at 62.5% for Fine-Grained Action), but performance declines to 32.3% for Event Attributes and 48.4% for Structural Event Logic. Furthermore, we note that while LLaVA-Video-72B and InternVL 3.5-38B achieve moderate performance on Events (32.4% and 33.3%), they perform only slightly above random chance in the other two categories.
  • Physical Dynamics. Performance is lowest in the Event Attributes category, which requires sensitivity to continuous or qualitative parameters such as kinematics (speed) and dynamics (force, magnitude). Strong proprietary models GPT-5 and Gemini 3 Pro only achieve 32.3% and 36.7% I-Acc, respectively, while the top-performing open-source model, Molmo2-8B, scores only 20.3%. Furthermore, several models, including Qwen3-VL-235B, LLaVA-Video-72B, and InternVL 3.5-38B, perform at near random chance, struggling to distinguish nuances such as gentle vs. forceful, fast vs. slow, or large vs. small amplitude. These results expose a systematic deficiency in current models’ understanding of low-level, physics-related temporal dynamics.
  • Gap in Structural Logic. We observe that GPT-5 and Gemini 3 Pro demonstrate relatively robust performance in Structural Logic Analysis, achieving 48.4% and 57.8%, respectively. However, a striking disparity emerges between these proprietary models and their open-source counterparts. The leading Qwen3-VL-235B achieves only 18.8%—trailing Gemini 3 Pro by 39.0%—and scores a mere 7.5% in the Causal Contingency sub-category. These results indicate that while open-source MLLMs may effectively detect isolated events, they lack the necessary logical abilities to reason about precise relationships between multiple events, particularly regarding causality.
Table 3: Category-Wise I-Acc (%) on TimeBlind. This table reports I-Acc for advanced models across 11 fine-grained temporal understanding tasks to pinpoint specific cognitive deficits. Due to space constraints, we use the following abbreviations for temporal categories: FG Action: Fine-Grained Action, State Trans: State Transitions, Mag: Magnitude, Dir: Direction, Dur: Duration, Rep: Repetition, Temp Topo: Temporal Topology, Causal Cont: Causal Contingency, and Cross Comp: Cross-Event Comparison. We include the overall performance (Avg) for each high-level category alongside the fine-grained categories. The results show clear performance gaps across categories. Models generally perform well on discrete Events, but struggle with Event Attributes, which require low-level physical understanding such as Speed and Force. The best results are bolded.
Models Event Event Attribute Structural Event Logic
[…]
5.3 Shortcut Analysis
To verify that solving TimeBlind requires genuine temporal understanding, we report three tests with GPT-5 in Table 4.
  • Single-Frame Bias. This experiment evaluates whether TimeBlind requires reasoning over a sequence rather than exploiting information from a single static frame. In this setting, the model is provided with the question and only one randomly sampled frame. The results show that GPT-5 performs poorly (4.5% I-Acc), indicating that TimeBlind requires sequential modeling.
    Table 4: Shortcut Analysis. We report I-Acc for three shortcut baselines using GPT-5 with 1 FPS sampling: Single Frame Bias (a question paired with a randomly selected video frame), Language Only (a question without visual input), and Visual-Cue (a question with shuffled video frames). The results demonstrate that solving TimeBlind requires genuine temporal understanding.
[…]
    Random Chance 50.0 25.0 25.0 6.3
    Single Frame 52.2 14.8 32.3 4.5
    Language Only 47.3 9.4 26.6 1.5
    Visual-Cue 49.6 10.7 23.2 3.0
  • Language-Only Bias. This experiment evaluates the influence of language priors within TimeBlind. In this setting, the model is provided with only the question without any visual information. As shown in Table 4, GPT-5 achieves only 1.5% I-Acc, demonstrating that visual information is essential for our benchmark setting.
  • Visual-Cue Shortcuts. Finally, we examine whether the benchmark can be solved by exploiting static visual cues that happen to correlate with temporal dynamics. In this setting, we shuffle the order of the input frames sampled at 1 FPS to ensure that successfully completing the task requires an understanding of the temporal sequence rather than merely detecting specific objects. GPT-5 achieves only 3.0% I-Acc in this setting, indicating that solving TimeBlind requires strict temporal understanding.
Moreover, Acc scores across all three settings hover around random chance, suggesting that correctly answering even a single question in TimeBlind requires strict temporal understanding. In summary, TimeBlind stands as a temporal-centric benchmark robust against shortcut solutions.
Table 5: Effect of Model Size and Input Frames. We report the I-Acc performance of several open-source and proprietary models across various model sizes and input frame counts. The results indicate that simply scaling up model size and frame count fails to significantly improve performance. Best results within each model grouping are highlighted in bold with a gray background.
Size Frames Q-Acc V-Acc Acc I-Acc
[…]
8B 8 34.7 41.5 61.9 17.1
16 35.6 42.7 62.6 18.5
32 35.6 43.1 62.9 17.4
14B 8 34.2 46.3 63.6 19.9
16 38.2 47.6 65.2 23.5
32 37.7 49.4 65.4 23.9
38B 8 35.8 42.8 63.2 20.4
16 38.9 45.4 64.5 24.4
32 39.4 46.4 65.9 25.1
[…]
7B 8 21.8 37.9 59.5 12.4
16 23.4 39.8 60.1 13.9
32 24.7 41.1 60.7 14.6
72B 8 29.7 46.2 63.5 19.9
16 32.4 48.4 65.0 23.0
32 32.8 48.7 65.2 23.3
[…]
8 60.9 69.2 79.0 49.1
16 62.7 69.9 79.5 50.9
32 61.1 68.2 78.7 48.3
Table 6: Inference-Time Reasoning. We compare standard Qwen3-VL and GPT-5 models against their reasoning-enabled Thinking counterparts across various settings, using the I-Acc metric at 1 FPS sampling. Results show that sufficient inference-time reasoning depth improves temporal understanding, but is far from sufficient to solve TimeBlind.
Model Mode I-Acc (%) $\Delta$
Qwen3-VL-8B Standard 19.6 -
[…]
Qwen3-VL-235B Standard 25.8 -
[…]
GPT-5 Standard 46.3 -
[…]
5.4 Additional Analysis
We conduct ablation studies on three factors: input frames, model size, and inference-time reasoning.
Number of Input Frames. Table 5.3 shows that increasing frames from 8 to 32 yields only marginal I-Acc gains: 1–5% for InternVL 3.5 and LLaVA-Video, and less than 2% for GPT-5. Even with sufficient frames, models struggle with TimeBlind, revealing fundamental limitations in fine-grained temporal understanding.
[…]
Test-Time Scaling (Reasoning). Table 6 shows inference-time reasoning results. Due to computational cost, we evaluate on a random 30% subset of the data. For Qwen3-VL, Thinking variants outperform Instruct models, with the 235B model gaining 10.4% but still achieving only 36.3% I-Acc. For GPT-5, varying reasoning effort from Low to High yields modest gains, peaking at 49.6% I-Acc—still far below human performance (98.2%). While deeper reasoning helps, it remains insufficient to solve TimeBlind.
6 Conclusion
We introduce TimeBlind, a diagnostic benchmark designed to rigorously assess the compositional spatio-temporal reasoning capabilities of MLLMs. By organizing evaluation around a structured cognitive taxonomy—spanning atomic events, event attribute, and structural event logic—and employing a strict minimal-pair design, we effectively isolate temporal understanding from static and linguistic shortcuts. Our evaluation reveals a significant gap: despite rapid progress in static vision-language tasks, current state-of-the-art models remain largely “time-blind,” trailing human performance by 50% and exhibiting fragility in fine-grained event attributes and temporal logic. We believe that TimeBlind will be beneficial for developing MLLMs capable of understanding fine-grained temporal dynamics.
Acknowledgments
This work was supported by Laboratory for Analytic Sciences via NC State University, ONR Award N00014-23-1-2356, NIH Award R01HD11107402, and Sony Focused Research award. We thank Zhiqiu Lin, Kewen Wu, and Deva Ramanan for their invaluable discussions during the development of this work.
Impact Statement
This work introduces a diagnostic benchmark for evaluating temporal understanding in video-language models. We anticipate several positive impacts: (1) TimeBlind can guide development of more temporally-aware models, which is critical for applications in robotics, autonomous driving, and assistive technologies; (2) by revealing systematic failure modes, our benchmark may help practitioners avoid deploying models in safety-critical scenarios where temporal reasoning is essential.
We also acknowledge potential concerns. Improved video understanding capabilities could enable more sophisticated surveillance or content moderation systems, raising privacy considerations. Additionally, our benchmark primarily features videos from controlled settings and internet sources, which may not represent the full diversity of real-world scenarios or populations. We encourage future work to expand temporal reasoning evaluation to more diverse contexts.
The benchmark itself does not enable direct harmful applications, as it evaluates rather than enhances model capabilities. We will release our data and code to support reproducible research.
Appendix A Appendix
Outline
This document provides additional details on TimeBlind and failure cases, supplementing the main paper. The outline is as follows:
  • Section A.1 The TimeBlind Benchmark
    • DATA GENERATION PROMPTS
    • BENCHMARK STATISTICS
    • TAXONOMY
  • Section LABEL:app_sec:failure_cases Failure Cases
A.1 The TimeBlind Benchmark
A.1.1 Data Generation Prompts
In this section, we present the full prompt template used in our Human–AI collaborative pipeline (see Section 3.3 in the main paper).
The system prompt for TimeBlind consists of three parts: (1) a General Prompt that defines the role and task, (2) Detailed Requirements that specify the core generation rules and logical constraints, and (3) an Output Format that standardizes the response structure. The prompt includes strict structural constraints to ensure high-quality minimal pairs. When using this prompt, we provide a single category as the “Input” each time.
Part 1: Role, Context, and Task Definition Role You are an expert in dataset curation and benchmark design for Multimodal Large Language Models (MLLMs). Context We are constructing a benchmark called “TimeBlind” to evaluate the temporal perception capabilities of MLLMs. In TimeBlind, each data instance consists of two videos ($v_{1},v_{2}$) and two questions ($q_{1},q_{2}$). The videos are visually similar (same objects/background) but contain distinct temporal dynamics. The questions are designed to distinguish these specific temporal features. During testing, the model receives only one video and one question at a time. Task I will provide you with a specific “Temporal Dynamic Category”. Your task is to generate a valid data instance (one pair of videos and one pair of questions) that perfectly fits the requirements below.
Part 2: Generation Requirements and Logical Constraints Requirements 1. Video Generation ($v_{1}$ and $v_{2}$) Format: Provide detailed text descriptions for both videos. Feasibility: The scenes must be easy to collect via web search, self-recording, or synthetic generation. Visual Consistency: $v_{1}$ and $v_{2}$ must share the same objects, background, and camera angle. They should look nearly identical in a static frame; the difference must be purely temporal. 2. Question Generation ($q_{1}$ and $q_{2}$) Type: Two-choice multiple-choice or Binary Questions (Yes/No). Relevance: Both questions must relate to the provided “Temporal Dynamic Category.” Temporal Dependency: The questions must strictly require temporal understanding. They cannot be solvable by looking at a single static frame. Logical Constraints (Crucial) The answers (AA) must adhere to the following logic matrix to ensure the model is truly perceiving time: Constraint A (Cross-Video Difference): The answer changes depending on the video. $A(q_{1},v_{1})\neq A(q_{1},v_{2})\qquad\text{and}\qquad A(q_{2},v_{1})\neq A(q_{2},v_{2})$ Constraint B (Intra-Video Difference): The two questions have different answers within the same video. $A(q_{1},v_{1})\neq A(q_{2},v_{1})\qquad\text{and}\qquad A(q_{1},v_{2})\neq A(q_{2},v_{2})$
Part 3: Output Specifications Output Format Please present the result in the following structured format: Temporal Category: [Input Category] Video Descriptions: $v_{1}$: [Description] $v_{2}$: [Description] Collection Source: [e.g., Synthetic / Self-shot / Web] Questions & Answers: $q_{1}$: [Question Text]
Answer in $v_{1}$: [Answer] / Answer in $v_{2}$: [Answer]
$q_{2}$: [Question Text]
Answer in $v_{1}$: [Answer] / Answer in $v_{2}$: [Answer]
Reason: Why single-frame is insufficient: [Brief explanation]
A.1.2 Benchmark Statistics
This section reports dataset statistics for TimeBlind. TimeBlind includes both two-choice multiple-choice and binary questions in equal numbers. It is a short-video benchmark, with an average video duration of 8.49 seconds.
Table 7: Statistics of the TimeBlind benchmark.
Statistic Value
Binary Question 1,200
Two-Choice Multiple-Choice Question 1,200
Avg. Video Length (sec) 8.49
Avg. Question Length (words) 20.54
A.1.3 Taxonomy
Table 8: Skill Definitions in TimeBlind for Video Event Understanding.
\__nicematrix_patch_booktabs:\__nicematrix_revert_colortbl:

### v2

@@
 Benchmark Size Event Types Event Attributes Temporal Topologies Causal Cross-Event Comparison Video Pairs Compl. Questions Human–Model Gap
-MVBench (li2024mvbench) 4K 2 2 2
-TOMATO (shangguan2024tomato) 1.4K 2 4 2 57.3%
-Vinoground (zhang2024vinoground) 2K 2 1 2 40.0%
-TempCompass (liu2024tempcompass) 4K 2 2 2
+MVBench (Li et al., 2024c) 4K 2 2 2
+TOMATO (Shangguan et al., 2024) 1.4K 2 4 2 57.3%
+Vinoground (Zhang et al., 2024a) 2K 2 1 2 40.0%
+TempCompass (Liu et al., 2024) 4K 2 2 2
 TimeBlind (Ours) 2.4K 2 6 13 50.0%
-Covers all 13 event relations in Allen’s Interval Algebra (allen1983maintaining): before, after, meets, met-by, overlaps, overlapped-by, starts, started-by, finishes, finished-by, during, contains, equals.
-Fine-grained spatio-temporal understanding is fundamental for long-horizon video reasoning (wang2025lvbench; pan2025basket) and embodied AI (tang2025video; yang2025cambrian).
+Covers all 13 event relations in Allen’s Interval Algebra (Allen, 1983): before, after, meets, met-by, overlaps, overlapped-by, starts, started-by, finishes, finished-by, during, contains, equals.
+Fine-grained spatio-temporal understanding is fundamental for long-horizon video reasoning (Wang et al., 2025b; Pan et al., 2025) and embodied AI (Tang et al., 2025; Yang et al., 2025b).
 Beyond simply recognizing what is present, an intelligent system must infer what changes, how it changes, and how multiple changes compose into causal structures.
-While recent Multimodal Large Language Models (MLLMs) (gemini; openai2025gpt5systemcard; bai2025qwen3vltechnicalreport; clark2025molmo2; cho2025perceptionlm) demonstrate impressive performance on general benchmarks, their “sense of time” remains surprisingly brittle.
+While recent Multimodal Large Language Models (MLLMs) (Pichai et al., 2025; OpenAI, 2025; Bai et al., 2025; Clark et al., 2025; Cho et al., 2025) demonstrate impressive performance on general benchmarks, their “sense of time” remains surprisingly brittle.
 As shown in Figure 1, even frontier models (e.g., GPT-5 and Gemini 3 Pro) struggle to distinguish atomic actions (e.g., shaking vs.
@@
 decelerating), and frequently fail to resolve linguistic temporal connectives (e.g., “as soon as”).
-This discrepancy suggests that current benchmark scores (jang2017tgif; activitynet; li2024mvbench; zhou2025mlvu; fu2025video; hu2025video) severely overestimate genuine temporal understanding capabilities.
+This discrepancy suggests that current benchmark scores (Jang et al., 2017a; Yu et al., 2019; Li et al., 2024c; Zhou et al., 2025a; Fu et al., 2025; Hu et al., 2025) severely overestimate genuine temporal understanding capabilities.
 The root issue lies in evaluation design.
-Existing benchmarks (activitynet; li2024mvbench; liu2024tempcompass; fu2025video) rarely isolate temporal structure as the sole discriminative factor.
-As a result, many models exploit “static shortcuts”—correlating visual entities with answers without modeling time (lei2023revealing; krojer2025shortcut).
-Furthermore, language priors (li2024naturalbench) allow models to guess answers based on textual plausibility.
+Existing benchmarks (Yu et al., 2019; Li et al., 2024c; Liu et al., 2024; Fu et al., 2025) rarely isolate temporal structure as the sole discriminative factor.
+As a result, many models exploit “static shortcuts”—correlating visual entities with answers without modeling time (Lei et al., 2023; Krojer et al., 2025).
+Furthermore, language priors (Li et al., 2024b) allow models to guess answers based on textual plausibility.
 To truly diagnose understanding, an evaluation must hold the visual content constant and vary only the temporal dynamics.
@@
 Unlike large-scale noisy benchmarks, TimeBlind prioritizes high-fidelity diagnostic precision.
-Similar to Winoground (thrush2022winoground), each instance serves as a rigorous test for a specific cognitive primitive, where high-quality annotations are prioritized over scale.
-Crucially, unlike prior work (zhang2024vinoground; krojer2025shortcut), TimeBlind evaluates logical composition, not just perception.
-Inspired by cognitive science, we extend the theory of Image Compositionality (krishna2017visual; li2024genai) to the temporal domain, organizing the benchmark around a hierarchical taxonomy: (i) Events, addressing what happened (e.g., changes in object attributes or action understanding); (ii) Event Attributes, describing how events unfold (e.g., speed, magnitude of change); and (iii) Structural Event Logic, examining how multiple events are composed (e.g., temporal topology, causality, and cross-event comparison).
+Similar to Winoground (Thrush et al., 2022), each instance serves as a rigorous test for a specific cognitive primitive, where high-quality annotations are prioritized over scale.
+Crucially, unlike prior work (Zhang et al., 2024a; Krojer et al., 2025), TimeBlind evaluates logical composition, not just perception.
+Inspired by cognitive science, we extend the theory of Image Compositionality (Krishna et al., 2017; Li et al., 2024a) to the temporal domain, organizing the benchmark around a hierarchical taxonomy: (i) Events, addressing what happened (e.g., changes in object attributes or action understanding); (ii) Event Attributes, describing how events unfold (e.g., speed, magnitude of change); and (iii) Structural Event Logic, examining how multiple events are composed (e.g., temporal topology, causality, and cross-event comparison).
 Within this framework, we encompass a diverse set of 11 fine-grained categories in real-world scenarios.
-For example, regarding Temporal Topology in Structural Event Logic, we incorporate Allen’s Interval Algebra (allen1983maintaining) by constructing videos that cover all 13 temporal event relations (e.g., overlaps, meets, equals) and designing questions that distinguish specific topology differences within each pair, going beyond simple sequencing relations (e.g., before, after) used in prior work.
-We evaluate over 20 state-of-the-art MLLMs, including leading proprietary models (e.g., Gemini 3 Pro (gemini), GPT-5 (openai2025gpt5systemcard)) as well as strong open-source models (e.g., Qwen3-VL (bai2025qwen3vltechnicalreport), Molmo2 (clark2025molmo2), PLM (cho2025perceptionlm)), on 600 curated instances (totaling 2,400 video-question pairs).
+For example, regarding Temporal Topology in Structural Event Logic, we incorporate Allen’s Interval Algebra (Allen, 1983) by constructing videos that cover all 13 temporal event relations (e.g., overlaps, meets, equals) and designing questions that distinguish specific topology differences within each pair, going beyond simple sequencing relations (e.g., before, after) used in prior work.
+We evaluate over 20 state-of-the-art MLLMs, including leading proprietary models (e.g., Gemini 3 Pro (Pichai et al., 2025), GPT-5 (OpenAI, 2025)) as well as strong open-source models (e.g., Qwen3-VL (Bai et al., 2025), Molmo2 (Clark et al., 2025), PLM (Cho et al., 2025)), on 600 curated instances (totaling 2,400 video-question pairs).
 We find that even the top-performing model, Gemini-3 Pro, achieves only 48.2% Instance Accuracy (I-Acc), which requires correctly distinguishing both videos in a pair, falling far below human performance (98.2%).
@@
 VideoQA Benchmarks.
-Early VideoQA datasets (activitynet; tgif; perception; xiao2021next; lei2018tvqa; xu2016msr; wu2024star) focus on simple scenarios with short clips and limited question types.
-Recent benchmarks address comprehensive evaluation (fu2025video; li2024mvbench; ma2025videoeval), complex reasoning (nagrani2025minerva; cheng2025video; hu2025video; song2025video), long-form understanding (mangalam2023egoschema; wang2025lvbench; wu2024longvideobench; song2024moviechat; zhou2025mlvu; yang2025egolife), and domain-specific settings (pan2025basket; long2025adsqa; yi2025exact).
-However, most do not isolate temporal structure as the sole discriminative factor, allowing models to exploit static shortcuts—relying on object co-occurrence or language priors without genuinely modeling temporal dynamics (lei2023revealing; krojer2025shortcut; li2024naturalbench).
+Early VideoQA datasets (Yu et al., 2019; Jang et al., 2017b; Patraucean et al., 2023; Xiao et al., 2021; Lei et al., 2018; Xu et al., 2016; Wu et al., 2024a) focus on simple scenarios with short clips and limited question types.
+Recent benchmarks address comprehensive evaluation (Fu et al., 2025; Li et al., 2024c; Ma et al., 2025), complex reasoning (Nagrani et al., 2025; Cheng et al., 2025; Hu et al., 2025; Song et al., 2025), long-form understanding (Mangalam et al., 2023; Wang et al., 2025b; Wu et al., 2024b; Song et al., 2024; Zhou et al., 2025a; Yang et al., 2025a), and domain-specific settings (Pan et al., 2025; Long et al., 2025; Yi et al., 2025).
+However, most do not isolate temporal structure as the sole discriminative factor, allowing models to exploit static shortcuts—relying on object co-occurrence or language priors without genuinely modeling temporal dynamics (Lei et al., 2023; Krojer et al., 2025; Li et al., 2024b).
 Evaluating Spatio-Temporal Compositionality.
 To assess the spatio-temporal understanding capabilities of MLLMs, several recent benchmarks have introduced diverse aspects of visual reasoning.
-In the image understanding domain, BLINK (fu2024blink) reformats and groups classic vision problems into perception-centric multiple-choice questions and carefully removes language priors.
-In the video understanding domain, several works explicitly target temporal understanding by designing temporally challenging questions (johnson2017clevr; shangguan2024tomato; cai2024temporalbench; xue2025seeing).
+In the image understanding domain, BLINK (Fu et al., 2024) reformats and groups classic vision problems into perception-centric multiple-choice questions and carefully removes language priors.
+In the video understanding domain, several works explicitly target temporal understanding by designing temporally challenging questions (Johnson et al., 2017; Shangguan et al., 2024; Cai et al., 2024; Xue et al., 2025).
 Beyond single-video evaluation, a growing line of work adopts paired video–question protocols to more precisely diagnose temporal reasoning.
-TempCompass (liu2024tempcompass) constructs paired videos by systematically manipulating an original video, such as reversing playback or altering temporal speed, thereby isolating temporal understanding from “static shortcuts.” VinoGround (zhang2024vinoground) further strengthens this paradigm by requiring models to answer identical questions over paired videos, where the correct answer is determined solely by temporal differences, effectively removing language priors.
-Follow-up works like GLIMPSE (zhou2025glimpse) and MVP (krojer2025shortcut) scale this approach to test physical and visual-centric reasoning.
+TempCompass (Liu et al., 2024) constructs paired videos by systematically manipulating an original video, such as reversing playback or altering temporal speed, thereby isolating temporal understanding from “static shortcuts.” VinoGround (Zhang et al., 2024a) further strengthens this paradigm by requiring models to answer identical questions over paired videos, where the correct answer is determined solely by temporal differences, effectively removing language priors.
+Follow-up works like GLIMPSE (Zhou et al., 2025b) and MVP (Krojer et al., 2025) scale this approach to test physical and visual-centric reasoning.
 Different from prior works, TimeBlind achieves spatio-temporal compositionality through a carefully curated formal taxonomy.
-Drawing inspiration from cognitive event perception (bach1986algebra) and extending the theory of image compositionality (krishna2017visual; thrush2022winoground; li2024naturalbench), we decompose temporal reasoning into atomic Events, parametric Event Attributes, and Structural Logic.
+Drawing inspiration from cognitive event perception (Bach, 1986) and extending the theory of image compositionality (Krishna et al., 2017; Thrush et al., 2022; Li et al., 2024b), we decompose temporal reasoning into atomic Events, parametric Event Attributes, and Structural Logic.
 3 The TimeBlind Benchmark
@@
 Diagnostic Metrics.
-Following rigorous diagnostic protocols (thrush2022winoground; li2024naturalbench), we report a hierarchy of metrics to mitigate shortcuts and assess the models’ genuine temporal understanding:
+Following rigorous diagnostic protocols (Thrush et al., 2022; Li et al., 2024b), we report a hierarchy of metrics to mitigate shortcuts and assess the models’ genuine temporal understanding:
   • Accuracy (Acc): Standard accuracy computed over all individual video–question trials.
@@
 III. Structural Event Logic. This level tests how multiple events relate and compose into higher-order structures.
-  • Temporal Topology: We adopt Allen’s Interval Algebra (allen1983maintaining) to comprehensively evaluate interval relations.
+  • Temporal Topology: We adopt Allen’s Interval Algebra (Allen, 1983) to comprehensively evaluate interval relations.
 Unlike prior work limited to simple sequencing (before/after), we cover all 13 relations in our videos, such as overlaps, meets, starts, and during.
@@
 Baseline Models.
-To thoroughly assess the challenges posed by TimeBlind, we evaluate over 20 MLLMs, including both frontier proprietary models (e.g., GPT-5 (openai2025gpt5systemcard), Gemini 3 Pro (gemini), Claude (anthropic_claude_sonnet_4_5_2025), and Qwen3-VL Plus (bai2025qwen3vltechnicalreport)) and state-of-the-art open-source models (e.g., Qwen3-VL (bai2025qwen3vltechnicalreport), InternVL 3.5 (internvl3), Molmo2 (clark2025molmo2), PLM (cho2025perceptionlm), Keye-VL (team2025kwai), MiniCPM-V-4.5 (yu2025minicpm), GLM-4.1V (hong2025glm), and Eagle 2.5 (chen2025eagle2.5)).
+To thoroughly assess the challenges posed by TimeBlind, we evaluate over 20 MLLMs, including both frontier proprietary models (e.g., GPT-5 (OpenAI, 2025), Gemini 3 Pro (Pichai et al., 2025), Claude (Anthropic, 2025), and Qwen3-VL Plus (Bai et al., 2025)) and state-of-the-art open-source models (e.g., Qwen3-VL (Bai et al., 2025), InternVL 3.5 (Wang et al., 2025a), Molmo2 (Clark et al., 2025), PLM (Cho et al., 2025), Keye-VL (Team et al., 2025), MiniCPM-V-4.5 (Yu et al., 2025), GLM-4.1V (Hong et al., 2025), and Eagle 2.5 (Chen et al., 2025)).
 Evaluation Metrics.
@@
 Best results are shown in bold.
-Model Size Q-Acc (%) V-Acc (%) Acc (%) I-Acc (%) $\uparrow$
+Model Size Q-Acc (%) V-Acc (%) Acc (%) I-Acc (%)
 Random Chance - 25.0 25.0 50.0 6.3
@@
 Model Size $\leq 10$B
-Keye-VL-1.5 (team2025kwai) 8B 17.5 26.5 55.6 6.6
-LLaVA-Video (zhang2024video) 7B 18.1 34.8 57.8 9.6
-InternVL3.5 (internvl3) 8B 30.9 39.0 59.3 13.3
-PLM-8B (cho2025perceptionlm) 8B 24.4 32.3 60.6 13.9
-Qwen3-VL (bai2025qwen3vltechnicalreport) 8B 34.5 43.6 64.8 19.7
-MiniCPM-V-4.5 (yu2025minicpm) 8B 31.7 43.6 62.7 18.3
-GLM-4.1V (hong2025glm) 9B 32.3 45.2 62.0 19.7
-Qwen3-VL (bai2025qwen3vltechnicalreport) 4B 29.9 43.7 62.6 17.7
-VideoChat-Flash (li2024videochat) 7B 31.7 45.4 64.9 19.7
-Eagle2.5 (chen2025eagle2.5) 8B 32.2 45.7 64.7 20.2
-Molmo2 (clark2025molmo2) 8B 41.0 52.4 68.7 31.2
+Keye-VL-1.5 (Team et al., 2025) 8B 17.5 26.5 55.6 6.6
+LLaVA-Video (Zhang et al., 2024b) 7B 18.1 34.8 57.8 9.6
+InternVL3.5 (Wang et al., 2025a) 8B 30.9 39.0 59.3 13.3
+PLM-8B (Cho et al., 2025) 8B 24.4 32.3 60.6 13.9
+Qwen3-VL (Bai et al., 2025) 8B 34.5 43.6 64.8 19.7
+MiniCPM-V-4.5 (Yu et al., 2025) 8B 31.7 43.6 62.7 18.3
+GLM-4.1V (Hong et al., 2025) 9B 32.3 45.2 62.0 19.7
+Qwen3-VL (Bai et al., 2025) 4B 29.9 43.7 62.6 17.7
+VideoChat-Flash (Li et al., 2024d) 7B 31.7 45.4 64.9 19.7
+Eagle2.5 (Chen et al., 2025) 8B 32.2 45.7 64.7 20.2
+Molmo2 (Clark et al., 2025) 8B 41.0 52.4 68.7 31.2
 Model Size $\geq 30$B
-InternVL 3.5 (internvl3) 30B 31.4 41.3 61.1 17.0
-LLaVA-Video (zhang2024video) 72B 28.8 45.0 63.2 18.5
-Qwen3-VL (bai2025qwen3vltechnicalreport) 235B 38.1 51.7 66.9 25.8
+InternVL 3.5 (Wang et al., 2025a) 30B 31.4 41.3 61.1 17.0
+LLaVA-Video (Zhang et al., 2024b) 72B 28.8 45.0 63.2 18.5
+Qwen3-VL (Bai et al., 2025) 235B 38.1 51.7 66.9 25.8
 […]
-Claude Sonnet 4.5 (anthropic_claude_sonnet_4_5_2025) - 30.6 38.5 59.2 13.6
-Qwen3-VL Plus (bai2025qwen3vltechnicalreport) - 38.7 52.6 67.3 26.0
-GPT-5 mini (openai2025gpt5systemcard) - 42.3 53.7 68.3 30.0
-Gemini 2.5 Flash (google_gemini_2_5_thinking_2025) - 47.6 53.7 68.9 33.2
-Gemini 2.5 Pro (google_gemini_2_5_thinking_2025) - 56.0 62.6 74.2 43.5
-GPT-5 (openai2025gpt5systemcard) - 58.5 67.2 77.3 46.3
-Gemini 3 Pro (gemini) - 60.2 66.0 76.2 48.2
+Claude Sonnet 4.5 (Anthropic, 2025) - 30.6 38.5 59.2 13.6
+Qwen3-VL Plus (Bai et al., 2025) - 38.7 52.6 67.3 26.0
+GPT-5 mini (OpenAI, 2025) - 42.3 53.7 68.3 30.0
+Gemini 2.5 Flash (Kavukcuoglu, 2025) - 47.6 53.7 68.9 33.2
+Gemini 2.5 Pro (Kavukcuoglu, 2025) - 56.0 62.6 74.2 43.5
+GPT-5 (OpenAI, 2025) - 58.5 67.2 77.3 46.3
+Gemini 3 Pro (Pichai et al., 2025) - 60.2 66.0 76.2 48.2
 In Table 5.1, we present the overall model performances on our benchmark, revealing that all models perform poorly in fine-grained temporal video understanding.
@@
 Outline
-This document provides additional details on TimeBlind and failure cases, supplementing the main paper. The outline is as follows:
+This document details TimeBlind and failure cases, supplementing the main paper. The outline is as follows:
   • Section A.1 The TimeBlind Benchmark
@@
 A.1.1 Data Generation Prompts
-In this section, we present the full prompt template used in our Human–AI collaborative pipeline (see Section 3.3 in the main paper).
+In this section, we present the full prompt template used in our Human–AI collaborative pipeline (see Section 3.3).
 The system prompt for TimeBlind consists of three parts: (1) a General Prompt that defines the role and task, (2) Detailed Requirements that specify the core generation rules and logical constraints, and (3) an Output Format that standardizes the response structure.
@@
 A.1.3 Taxonomy
-Table 8: Skill Definitions in TimeBlind for Video Event Understanding.
+Table 8: Hierarchical taxonomy of TimeBlind for video understanding.
+The taxonomy is structured into three levels—Event, Event Attribute, and Structural Event Logic.
+For each level, we present its conceptual definition, associated sub-categories, and representative paired-video descriptions for each sub-category.
 \__nicematrix_patch_booktabs:\__nicematrix_revert_colortbl: